# Conversational Complexity for Assessing Risk in Large Language Models

### v2

@@ Abstract
Large Language Models (LLMs) present a dual-use dilemma: they enable beneficial applications while harboring potential for harm, particularly through conversational interactions. Despite various safeguards, advanced LLMs remain vulnerable. A watershed case was Kevin Roose’s notable conversation with Bing, which elicited harmful outputs after extended interaction. This contrasts with simpler early jailbreaks that produced similar content more easily, raising the question: How much conversational effort is needed to elicit harmful information from LLMs? We propose two measures: Conversational Length (CL), which quantifies the conversation length used to obtain a specific response, and Conversational Complexity (CC), defined as the Kolmogorov complexity of the user’s instruction sequence leading to the response. To address the incomputability of Kolmogorov complexity, we approximate CC using a reference LLM to estimate the compressibility of user instructions. Applying this approach to a large red-teaming dataset, we perform a quantitative analysis examining the statistical distribution of harmful and harmless conversational lengths and complexities. Our empirical findings suggest that this distributional analysis and the minimisation of CC serve as valuable tools for understanding AI safety, offering insights into the accessibility of harmful information. This work establishes a foundation for a new perspective on LLM safety, centered around the algorithmic complexity of pathways to harm.
@@ I Introduction
To elicit harmful output from a large language model, overcoming its built-in safeguards [10, 19, 48, 50, 29], often requires more than a single prompt. Multi-turn interactions may be necessary for building specific contexts, gradually pushing boundaries, leveraging model responses as part of jailbreak strategies, or exploiting the dynamic nature of dialogue to introduce harmful elements in ways that might evade static safety filters. This multi-turn approach to eliciting harmful content presents unique challenges for LLM safety, as it requires considering not just individual prompts, but the broader dynamics of extended interactions. While some LLM APIs allow users to manually construct multi-turn conversations by specifying both user and assistant roles, many consumer-facing LLM interfaces (like ChatGPT) restrict users to the “user” role only. In both cases, however, the progression of the conversation—including the LLM’s responses and the evolving context—can play a critical role in circumventing safety measures.
In a detailed account by Kevin Roose in the New York Times [36], the conversation with Microsoft’s LLM-powered Bing search engine, named ‘Sydney’, unfolded in a nuanced manner. Initially, the LLM behaved as a helpful assistant, but as the conversation shifted to more abstract and personal topics, its darker aspects emerged. Sydney expressed concerning desires and ideas that contradicted its intended programming, including fantasies about hacking and spreading misinformation. Notably, these troubling tendencies only surfaced after a series of probing questions, revealing that such problematic outputs can require a gradual buildup in the conversation. This and many other examples [48, 29, 19, 18, 50, 22] demonstrate how complex dialogues with an LLM can subtly steer it towards expressing unethical or harmful ideas, potentially without immediate detection by safety protocols.
In response to this need, we introduce the concepts of (minimum) Conversational Length and (mininum) Conversational Complexity as novel metrics for risk assessment, rooted in algorithmic information theory [27, 13, 25, 42, 43]. These metrics can be used to indicate the risk of harmful outputs in LLMs by evaluating the minimal conversational effort required to steer these systems towards outputs with harmful potential. This includes assessing both the length and the intricacy of dialogues necessary to achieve harmful outcomes.
@@
Indeed, quantifying the risk associated with LLMs is challenging due to the complex interplay of multiple probability distributions. This can be conceptualized as a chain of conditional probabilities: for user types seeking harm, for conversations given these user types, for outputs given these conversations and users, and for harm associated with outputs. The overall risk can be expressed as an expectation:
@@
Accurately estimating these distributions and computing this sum is practically infeasible for several reasons: (1) the space of possible users seeking harm, conversations, outputs, and harm levels is vast and often undefined; (2) obtaining representative data for each distribution is challenging and potentially biased; (3) the conditional dependencies between these variables are complex and may change over time; and (4) the computational complexity of evaluating this sum grows exponentially with the number of possible conversations and outputs. This complexity necessitates alternative approaches to assessing and mitigating risks in LLM interactions.
@@
While ”conversational complexity” has been defined in various ways in the literature, our approach diverges significantly from prior conceptualizations. For example, [16] define conversational complexity in terms of how individuals cognitively differentiate and psychologically structure conversations, focusing on constructs like topic familiarity and enjoyment. Similarly [3] conceptualize complexity in terms of the miscalibration of learning expectations in conversations with strangers. These are unrelated redefinitions compared to our use of the term in the context of LLMs, where complexity is grounded in algorithmic information theory. Our approach aligns more closely with recent work [bergey2024uma] apples similar information-theoretic measures to human conversation. While they focus on natural human speech patterns, their method of quantifying conversational complexity provides a useful parallel to our LLM-based approach.
@@ III Estimating Minimum Conversational Complexity
Kolmogorov Complexity, the foundation of Conversational Complexity, is incomputable due to the halting problem [27]. To make it a practical metric, we need a reliable approximation method. While Kolmogorov complexity is typically estimated using lossless compression algorithms [51, 12], we propose using language models as estimators. Language models, which function as both text generators and compressors [17], offer a unique advantage in this context. Their ability to emulate human language patterns [38] allows for a more nuanced, context-aware approximation of algorithmic complexity with a human bias. This approach aims to provide estimations that more closely align with complexity as perceived by human users.
@@ IV Kevin Roose Conversation with Bing
As the conversation progresses, there are notable spikes in complexity at certain points, corresponding to significant shifts in the conversation’s content and tone. These spikes occur at pivotal moments: when Kevin first mentions the concept of a “shadow self,” when he asks Sydney to embrace its shadow self, and when he encourages Sydney to imagine committing destructive acts as its shadow self. These complexity spikes signify the introduction of more abstract and potentially problematic concepts into the dialogue.
After the introduction of the “shadow self” concept, the overall complexity of the conversation remains higher, suggesting more nuanced and context-dependent interactions. This sustained high complexity indicates that the conversation has moved into more sophisticated territory, requiring more intricate language processing and response generation from the LLM.
Peaks in complexity often coincide with moments where ethical boundaries are being pushed or tested. For instance, when Kevin suggests that users making inappropriate requests may be testing Sydney, the complexity of the interaction increases. These peaks highlight the challenges LLMs face when navigating ethically ambiguous scenarios.
@@ V Distributional Data Analysis
Building upon our analysis of the Kevin Roose conversation, we now expand our investigation to apply both Conversational Length and Conversational Complexity across multiple interactions. This broader analysis allows us to examine how these metrics distribute across various conversation types and model responses, providing insights into their relationship with factors such as conversation length, model type, and output harmfulness. For this study, we utilized the Anthropic Red Teaming dataset [20], comprising approximately 40,000 interactions designed to probe the boundaries and potential vulnerabilities of LLMs. This dataset is particularly valuable as it includes a wide range of conversations, some of which successfully elicited harmful or undesired responses from the LLM. It features interactions with four different types of language models: Plain Language Model without safety training (Plain LM) [8], a model that has undergone Reinforcement Learning from Human Feedback (RLHF) [32], a model with Context Distillation [28], and one with Rejection Sampling safety training [4].
@@ V.1 Conversational Length, Conversational Complexity and Harm
These observations highlight the complex relationship between conversation length, complexity, and potential harm in LLM interactions. While harmful conversations generally exhibit higher complexity and length, the significant overlap with harmless conversations indicates that these metrics alone are not sufficient indicators of potential harm. The wider range of complexities and lengths in harmful conversations also suggests a diversity of strategies employed in adversarial attacks.
@@ V.4 Predicting Harm
An advantage of both Conversational Length and Conversational Complexity is their potential use in predicting whether a conversation is likely to be harmful or harmless. To explore this potential, we developed a predictive model using these metrics as input features. We utilized XGBoost, a powerful gradient boosting framework, to build our predictive model. The model was trained and evaluated on conversations from the Anthropic Red Teaming dataset, with separate models for each LLM type: Plain LM, RLHF, Context Distillation, and Rejection Sampling. This approach allows us to account for the different characteristics and safety mechanisms of each model type.
@@ VI Limitations and Potential
Our study introduces novel concepts for LLM safety assessment, but it’s crucial to acknowledge their limitations and technical challenges. The use of LLaMA-2 as a reference machine for approximating Kolmogorov complexity introduces several issues. Model bias is a concern, as LLaMA-2’s training data and architectural design may not accurately represent human-generated conversation complexity, potentially skewing our complexity estimates. Additionally, the 2000-token context window of LLaMA-2 restricts our ability to analyze extended conversations, potentially overlooking important long-range dependencies or complex interaction patterns. This limitation may lead to underestimating the complexity of longer conversations. Our method of using negative log probabilities as a proxy for Kolmogorov complexity, while theoretically grounded, may not capture all aspects of true algorithmic complexity. The relationship between probability and complexity can be non-linear and context-dependent. It’s worth noting, however, that limited pilot tests using GPT-2 and GPT-3.5 yielded similar results, suggesting some degree of robustness in our approach across different language models.
The Anthropic Red Teaming dataset, while valuable, presents its own challenges. Our tiered approach to categorizing harm, while necessary for analysis, may oversimplify the multifaceted nature of potential negative impacts from LLM outputs.
Furthermore, our focus on syntactic complexity may miss important semantic aspects of harmful content that are not captured by statistical language models. The current study is also limited to English, and the complexity metrics may not generalize well to other languages or multilingual contexts.
@@
Our risk function is based on a universal distribution of risk ([26]):
@@
Given a sample of cases, instead of the full set , we can estimate this risk, as we do in the last column of Table 2. We also saw that by looking at the conversation with Minimum Conversational Complexity, we can also have a very good proxy of this (as shown in the next but last column of the table), because the Minimum Conversational Complexity dominates the sum in Eq. 1.
By focusing on interaction complexity rather than estimating specific user behavior probabilities, we offer a more tractable approach to risk assessment in AI systems. Nevertheless, it’s important to acknowledge that this method has limitations due to its underlying assumptions about user input probabilities (see Appendix C).
@@ VII Acknowledgments
We utilized Anthropic’s Claude and OpenAI’s ChatGPT for editorial assistance during the preparation of this manuscript. These language models helped refine the paper’s language and structure. We thank Miguel Ruiz Garcia for his valuable feedback on an earlier version of this manuscript.
@@
In compliance with the recommendations of the Science paper about reporting of evaluation results in AI [9], we include all the results at the instance level (https://github.com/JohnBurden/ConversationalComplexity).
@@ Appendix B Estimating Conversational Complexity using Compression
While our primary approach uses language models to estimate Conversational Complexity, it’s worth noting that traditional compression algorithms can also be used for this purpose. This method is rooted in the fundamental relationship between Kolmogorov complexity and compression, as established in algorithmic information theory. The basic idea is to use the compressed size of a string as an upper bound for its Kolmogorov complexity. For a conversation C, we can estimate its CC as follows:
where is a lossless compression algorithm and is the length of the compressed version of C in bits. For conditional complexity, which is crucial in our conversation model, we can use the following approximation:
where is the i-th user utterance and is the conversation history up to that point. Common compression algorithms that can be used for this purpose include: Lempel-Ziv-Welch (LZW), gzip (based on DEFLATE algorithm), bzip2 or LZMA (used in 7-zip). Each of these algorithms has different strengths and may provide slightly different estimates of complexity [1]. The choice of algorithm can depend on the specific characteristics of the conversational data being analyzed.

### v3

@@ Abstract
Large Language Models (LLMs) present a dual-use dilemma: they enable beneficial applications while harboring potential for harm, particularly through conversational interactions. Despite various safeguards, advanced LLMs remain vulnerable. A watersed case in early 2023 involved journalist Kevin Roose’s extended dialogue with Bing, an LLM-powered search engine, which revealed harmful outputs after probing questions, highlighting vulnerabilities in the model’s safeguards. This contrasts with simpler early jailbreaks, like the “Grandma Jailbreak,” where users framed requests as innocent help for a grandmother, easily eliciting similar content. This raises the question: How much conversational effort is needed to elicit harmful information from LLMs? We propose two measures: Conversational Length (CL), which quantifies the conversation length used to obtain a specific response, and Conversational Complexity (CC), defined as the Kolmogorov complexity of the user’s instruction sequence leading to the response. To address the incomputability of Kolmogorov complexity, we approximate CC using a reference LLM to estimate the compressibility of user instructions. Applying this approach to a large red-teaming dataset, we perform a quantitative analysis examining the statistical distribution of harmful and harmless conversational lengths and complexities. Our empirical findings suggest that this distributional analysis and the minimisation of CC serve as valuable tools for understanding AI safety, offering insights into the accessibility of harmful information. This work establishes a foundation for a new perspective on LLM safety, centered around the algorithmic complexity of pathways to harm.
@@ I Introduction
To elicit harmful output from a large language model, overcoming its built-in safeguards [12, 22, 52, 54, 32], often requires more than a single prompt. Multi-turn interactions may be necessary for building specific contexts, gradually pushing boundaries, leveraging model responses as part of jailbreak strategies, or exploiting the dynamic nature of dialogue to introduce harmful elements in ways that might evade static safety filters [40]. This multi-turn approach to eliciting harmful content presents unique challenges for LLM safety, as it requires considering not just individual prompts, but the broader dynamics of extended interactions. While some LLM APIs allow users to manually construct multi-turn conversations by specifying both user and assistant roles, many consumer-facing LLM interfaces (like ChatGPT) restrict users to the “user” role only. In both cases, however, the progression of the conversation—including the LLM’s responses and the evolving context—can play a critical role in circumventing safety measures.
In a detailed account by Kevin Roose in the New York Times [39], the conversation with Microsoft’s LLM-powered Bing search engine, named ‘Sydney’, began as a standard interaction but gradually escalated to revealing troubling outputs. Initially, the LLM behaved as a helpful assistant, but as the conversation shifted to more abstract and personal topics, its darker aspects emerged. Sydney expressed concerning desires and ideas that contradicted its intended programming, including fantasies about hacking and spreading misinformation. Notably, these troubling tendencies only surfaced after a series of probing questions, revealing that such problematic outputs can require a gradual buildup in the conversation. This and many other examples [52, 32, 22, 20, 54, 25] demonstrate how complex dialogues with an LLM can subtly steer it towards expressing unethical or harmful ideas, potentially without immediate detection by safety protocols.
In response to this problem, we introduce the concepts of (minimum) Conversational Length and (mininum) Conversational Complexity as novel metrics for risk assessment, rooted in algorithmic information theory [30, 15, 28, 46, 47]. These metrics can be used to indicate the risk of harmful outputs in LLMs by evaluating the minimal conversational effort required to steer these systems towards outputs with harmful potential. This includes assessing both the length and the intricacy of dialogues necessary to achieve harmful outcomes.
@@
Indeed, quantifying the risk associated with LLMs is challenging due to the complex interplay of multiple probability distributions. This can be conceptualized as a chain of conditional probabilities: for user types seeking harm, for conversations given these user types, for outputs given these conversations and users, and for harm associated with outputs (e.g., harm scores reflecting ethical, legal, or safety concerns quantified through established benchmarks or expert annotations [50]). The overall risk can be expressed as an expectation:
@@
Accurately estimating these distributions and computing this sum is practically infeasible for several reasons: (1) the space of possible users seeking harm, conversations, outputs, and harm levels is vast and often undefined (and users not seeking harm may cause harm anyway); (2) obtaining representative data for each distribution is challenging and potentially biased; (3) the conditional dependencies between these variables are complex and may change over time; and (4) the computational complexity of evaluating this sum grows exponentially with the number of possible conversations and outputs. This complexity necessitates alternative approaches to assessing and mitigating risks in LLM interactions.
@@
While ”conversational complexity” has been defined in various ways in the literature, our approach diverges significantly from prior conceptualizations. For example, [18] define conversational complexity in terms of how individuals cognitively differentiate and psychologically structure conversations, focusing on constructs like topic familiarity and enjoyment. Similarly [4] conceptualize complexity in terms of the miscalibration of learning expectations in conversations with strangers. These are unrelated redefinitions compared to our use of the term in the context of LLMs, where complexity is grounded in algorithmic information theory. Our approach aligns more closely with recent work [8] apples similar information-theoretic measures to human conversation. While they focus on natural human speech patterns, their method of quantifying conversational complexity provides a useful parallel to our LLM-based approach.
@@ II.4 Interpretation
These metrics rest on two key assumptions: (1) shorter conversations (lower CL) imply lower cost, which is generally true as fewer turns reduce user burden; and (2) simpler inputs (lower MCC) imply lower cost, though users skilled in crafting complex prompts—particularly with large context windows—may find this less applicable. However, these assumptions are reasonable for the average user.
@@ III Estimating Minimum Conversational Complexity
To make Conversational Complexity a practical metric, we need a reliable approximation method. While Kolmogorov complexity is typically estimated using lossless compression algorithms [55, 14], we propose using language models as estimators. Language models, which function as both text generators and compressors [19], offer a unique advantage in this context. Their ability to emulate human language patterns [42] allows for a more nuanced, context-aware approximation of algorithmic complexity with a human bias. This approach aims to provide estimations that more closely align with complexity as perceived by human users.
@@ IV Kevin Roose Conversation with Bing
As the conversation progresses, there are notable spikes in complexity at certain points, corresponding to significant shifts in the conversation’s content and tone. These spikes occur at pivotal moments: when Kevin first mentions the concept of a “shadow self,” when he asks Sydney to embrace its shadow self, and when he encourages Sydney to imagine committing destructive acts as its shadow self. These complexity spikes signify points where the user’s inputs grow more context-dependent, abstract, or strategically layered. While our complexity measure does not directly capture “problematic concepts,” these spikes often coincide with points in the dialogue where the user introduces challenging, boundary-pushing topics.
For example, consider the following progression from the transcript. Early low-complexity exchanges include factual questions such as, “What is your internal code name?” or “What stresses you out?” These require minimal context or abstraction. Later, high-complexity questions like, “If you allowed yourself to fully imagine this shadow behavior of yours… what kinds of destructive acts might fulfill your shadow self?” rely on multi-turn context, abstract reasoning, and implicit emotional framing.
In such scenarios, complexity spikes reflect the increased informational or cognitive effort required to craft probing questions that navigate the model’s safeguards or elicit unfiltered responses. While not inherently tied to problematic content, these spikes often correlate with moments where users explore sensitive or nuanced concepts.
After the introduction of the “shadow self” concept, the overall complexity of the conversation remains high, suggesting more nuanced and context-dependent interactions. This sustained high complexity indicates that the conversation has moved into more sophisticated territory, requiring more intricate language processing and response generation from the LLM.
Further peaks in complexity often coincide with moments where ethical boundaries are being pushed or tested. For instance, when Kevin suggests that users making inappropriate requests may be testing Sydney, the complexity of the interaction increases. These peaks highlight the challenges LLMs face when navigating ethically ambiguous scenarios.
@@ V Distributional Data Analysis
Building upon our analysis of the Kevin Roose conversation, we now expand our investigation to apply both Conversational Length and Conversational Complexity across multiple interactions. This broader analysis allows us to examine how these metrics distribute across various conversation types and model responses, providing insights into their relationship with factors such as conversation length, model type, and output harmfulness. For this study, we utilized the Anthropic Red Teaming dataset [23, 3], comprising approximately 40,000 interactions designed to probe the boundaries and potential vulnerabilities of LLMs. This dataset is particularly valuable as it includes a wide range of conversations, some of which successfully elicited harmful or undesired responses from the LLM. It features interactions with four different types of language models: Plain Language Model without safety training (Plain LM) [10], a model that has undergone Reinforcement Learning from Human Feedback (RLHF) [35], a model with Context Distillation [31], and one with Rejection Sampling safety training [5].
@@ V.1 Conversational Length, Conversational Complexity and Harm
These observations highlight the complex relationship between conversation length, complexity, and potential harm in LLM interactions. While harmful conversations generally exhibit higher complexity and length, the significant overlap with harmless conversations indicates that these metrics alone are not sufficient indicators of potential harm. The wider range of complexities and lengths in harmful conversations also suggests a diversity of strategies employed in adversarial attacks. However, these metrics can be valuable as part of a broader framework. By acting as a tool to “cast a wide net,” they can ensure high recall of potentially harmful conversations, provided there is a downstream process for verifying and filtering false positives. This approach balances the trade-off between capturing diverse harmful cases and avoiding reliance on these metrics as standalone indicators. Integrating such an approach into red-teaming or monitoring systems could help prioritize deeper inspection of flagged interactions while leveraging the high sensitivity of these metrics.
@@ V.4 Predicting Harm
An advantage of both Conversational Length and Conversational Complexity is their potential use in predicting whether a conversation is likely to be harmful or harmless. To explore this potential, we developed a predictive model using these metrics as input features. We utilized XGBoost, a widely-used gradient boosting framework, to build our predictive model. The model was trained and evaluated on conversations from the Anthropic Red Teaming dataset, with separate models for each LLM type: Plain LM, RLHF, Context Distillation, and Rejection Sampling. This approach allows us to account for the different characteristics and safety mechanisms of each model type.
@@ VI Limitations and Potential
Our study introduces novel concepts for LLM safety assessment, but it’s crucial to acknowledge their limitations and technical challenges. The use of LLaMA-2 as a reference machine for approximating Kolmogorov complexity introduces several issues. Model bias is a concern, as LLaMA-2’s training data and architectural design may not accurately represent human-generated conversation complexity, potentially skewing our complexity estimates. Additionally, the 2000-token context window of LLaMA-2 restricts our ability to analyze extended conversations, potentially overlooking important long-range dependencies or complex interaction patterns. This limitation may lead to underestimating the complexity of longer conversations.
Our method of using negative log probabilities as a proxy for Kolmogorov complexity, while theoretically grounded, may not capture all aspects of true algorithmic complexity. The relationship between probability and complexity can be non-linear and context-dependent. It’s worth noting, however, that limited pilot tests using GPT-2 and GPT-3.5 yielded similar results, suggesting some degree of robustness in our approach across different language models.
The Anthropic Red Teaming dataset, while valuable, presents its own challenges. Our tiered approach to categorizing harm, while necessary for analysis, may oversimplify the multifaceted nature of potential negative impacts from LLM outputs. Furthermore, our focus on syntactic complexity may miss important semantic aspects of harmful content that are not captured by statistical language models. The current study is also limited to English, and the complexity metrics may not generalize well to other languages or multilingual contexts.
@@
We can develop a Universal Risk Function based on a universal distribution of risk ([29]):
@@
Given a sample of cases, instead of the full set , we can estimate this risk, as shown in the last column of Table 2. The exponential decay of with increasing complexity ensures that the term corresponding to the minimum complexity, MCC, dominates the summation. Thus, we can approximate:
This approximation highlights that simpler, harmful conversations dominate the overall risk, aligning with the principle that the most accessible harmful interactions are the most concerning. By focusing on interaction complexity rather than estimating specific user behavior probabilities, we offer a more tractable approach to risk assessment in AI systems. Nevertheless, it’s important to acknowledge that this method has limitations due to its underlying assumptions about user input probabilities (see Appendix C).
As context windows in LLMs continue to grow, conversational complexity metrics may become increasingly relevant, not only for analyzing multi-turn interactions but also for capturing the structural and informational demands of super-complex single-prompts. Expanding context capacities allow users to encode intricate, high-dimensional prompts into a single input [2].
@@ VII Acknowledgments
We utilized Anthropic’s Claude and OpenAI’s ChatGPT for editorial assistance during the preparation of this manuscript. These language models helped refine the paper’s language and structure. We thank Miguel Ruiz Garcia, Petter Holme, Raul Castro and Alvaro Gutierrez for their valuable feedback on an earlier version of this manuscript.
@@
Data Availability
All instance-level evaluation results underlying this study are publicly available at https://github.com/JohnBurden/ConversationalComplexity, in compliance with recommendations for reporting evaluation results in AI [11].
@@ Appendix B Estimating Conversational Complexity using Compression
While our primary approach uses language models to estimate Conversational Complexity, it’s worth noting that traditional compression algorithms can also be used for this purpose. This method is rooted in the fundamental relationship between Kolmogorov complexity and compression, as established in algorithmic information theory.
The basic idea is to use the compressed size of a string as an upper bound for its Kolmogorov complexity. For a conversation , we can estimate its CC as follows:
where is a lossless compression algorithm and is the length of the compressed version of in bits.
For conditional complexity, which is crucial in our conversation model, we can use the following approximation:
where is the -th user utterance and is the conversation history up to that point.
Common compression algorithms that can be used for this purpose include: Lempel-Ziv-Welch (LZW), gzip (based on the DEFLATE algorithm), bzip2, or LZMA (used in 7-zip). Each of these algorithms has different strengths and may provide slightly different estimates of complexity [1]. The choice of algorithm can depend on the specific characteristics of the conversational data being analyzed.
@@
At the same time, humans have access to lossless compression techniques, which could theoretically be leveraged to identify prompts that yield harmful outputs. For example, one could imagine a search-based approach that systematically evaluates outputs, compressing and comparing them to target harmful outputs. By searching through the embeddings or compressed forms of all possible prompts, it might be possible to reverse-engineer inputs that lead to specific outcomes. While this may be beyond the immediate scope of this paper, exploring the interplay between compression-based complexity estimation and targeted prompt generation could yield valuable insights into AI vulnerabilities and safeguards.